OpenStudy (anonymous):

For the following system of equations, what is the x-value of the solution? -x+2y=6 6y= x+18 a. -18 b. 0 c. -6 d. 3

OpenStudy (zale101):

Use the substitution method. Solve for one variable from one equation and plug the equation of the same variable into the other equation.

OpenStudy (zale101):

Clear?

OpenStudy (zale101):

-x+2y=6 <--i'll choose this equation and i will solve for y 6y= x+18 \(\Large -x+2y=6\) add x to both sides \(\Large -x+x+2y=6+x\) \(\Large 2y=6+x\) divide everything by 2 \(\Large(\frac{1}{2}) 2y=(\frac{1}{2})(6+x)\) \(\Large y=3+ \frac{1}{2}x\) Plug what we have for the equation for y into the y in this equation 6y= x+18

OpenStudy (anonymous):

So b then. I got 2x=0 so it would be x=0. Right?

OpenStudy (anonymous):

I did 0. it was right. Thank you!